OpenStudy (anonymous):

Calculate the discriminant of 5x^2+10x-15=0 and then describe the type of solution you will get?

OpenStudy (anonymous):

^? 5x^(2)+10x-15=0 Factor out the GCF of 5 from each term in the polynomial. 5(x^(2))+5(2x)+5(-3)=0 Factor out the GCF of 5 from 5x^(2)+10x-15. 5(x^(2)+2x-3)=0 In this problem 3*-1=-3 and 3-1=2, so insert 3 as the right hand term of one factor and -1 as the right-hand term of the other factor. 5(x+3)(x-1)=0 Divide both sides of the equation by 5. Dividing 0 by any non-zero number is 0. (x+3)(x-1)=0 Set each of the factors of the left-hand side of the equation equal to 0. x+3=0_x-1=0 Since 3 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3 from both sides. x=-3_x-1=0 Set each of the factors of the left-hand side of the equation equal to 0. x=-3_x-1=0 Since -1 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 1 to both sides. x=-3_x=1 The complete solution is the set of the individual solutions. x=-3,1